OpenStudy (anonymous):

There is a following example in one of the 18.01 lectures: y^4 + xy^2 - 2 = 0 y' = -y^2 / (4y^3 + 2xy) What is the reason that y does not get cancelled?

sam (.sam.):

\[4y^3\frac{dy}{dx}+(2xy\frac{dy}{dx}+y^2)=0 \] \[\frac{dy}{dx}(4y^3+2xy)+y^2=0\] \[y'=\frac{-y^2}{4y^3+2xy}\]

OpenStudy (anonymous):

Ye, hence the question why we are not cancelling one of the "y" in the right hand side.

sam (.sam.):

Oh you can cancel

sam (.sam.):

there's nothing wrong but leaving this form is the same

OpenStudy (anonymous):

because if they canel and when intergrating it,it is not going to give you the original equestion

sam (.sam.):

Its the same

OpenStudy (anonymous):

Thanks, that's what I thought initially. Just felt weird that no one asked this at the lecture, like I was missing something obvious.